OpenStudy (anonymous):

HELP!! Will medal! What is the simplified form> 3a^4b^-2c^3

OpenStudy (anonymous):

Can you write that out using the equation editor please?

OpenStudy (anonymous):

\[3a ^{4}b ^{-2}c ^{3}\] @Indivicivet

OpenStudy (anonymous):

I don't think there is a simplified version of that! You could possibly bring the b² down to the bottom, but I don't know if I'd call that simplifying. I can only assume that's what is expected.

OpenStudy (anonymous):

@Indivicivet, This is what I have for answers. \[\frac{ 81a^4 }{ b^2c^3 }\]\[\fr\[\frac{ 3a^4 }{ b^2c^3 }\]ac{ 81a^4c^3 }{ b^2 }\]\[\frac{ 3a^4c^3 }{ b^2}\]

OpenStudy (anonymous):

The 4th one was the rearrangement I was talking about ^_^ Can you see why the others are not equivalent? Also, I disapprove of this question; I still don't think that's worth being called simplifying ^^;

OpenStudy (anonymous):

Nither do I. All you do is put it into fraction form.